# spacing=nonfrench

### v1

@@ 3.4 Ownership and Non-Blocking Communication
Using the data buffer concept instead of raw pointers already offers benefits such as an ergonomic interaction with standard library containers, improved type safety and the potential to prevent out-of-bounds accesses. Another important advantage of this object-oriented data handling is that it allows us to model a proper concept of ownership. In C++, if an object owns a resource—such as heap-allocated memory—the object is responsible for releasing this resource when it goes out of scope. This is usually achieved by placing the clean-up code in the object’s destructor and the key idea behind the RAII idiom. A data buffer owning its underlying memory is therefore responsible for freeing this memory once it goes out of scope, thus preventing memory leaks. This plays well with C++’s move semantics. During move-construction or move-assignment, ownership of the underlying resource is transferred from the source to the destination object. Using this mechanism it is for example possible to move a (receive) data buffer to an MPI call as shown in Fig. 6. Then the MPI call receives data into the provided buffer, which it now owns. Once the communication is completed, it returns the updated receive buffer by value resulting in a clean and idiomatic pass through of the receive data memory. This data flow allows us to avoid the use of out parameters that are discouraged in C++ [25, F.20].
In terms of library interface design this “pass through” of data can be achieved by considering the value category of data buffers passed to communication calls: If a data buffer is an r-value, the data buffer is moved to the MPI operation, and returned to the caller as a return value of the function call. If a data buffer is an l-value, it is passed by reference to the operation, ownership is not transferred, and it is not returned. For an example see Fig. 6.
@@
A robust solution to this problem is to transfer ownership of data buffers involved in non-blocking communication calls to the C++ interface, which forward it to request object as show in Fig. 6. The request object is conceptually similar to std::future  [^3] Using std::future to provide a safe interface for non-blocking communicating is not possible, as they are tied to asynchronous progress happening in the background, which the MPI standard does not guarantee.: Users can only access the data upon completing the request via wait or test, which move the data back to the caller, either directly by value in the former case, or encapsulated in std::optional in the latter. This way, invalid accesses to data involved in non-blocking communication can be prevent entirely through library semantics. This idea was first introduced and implemented in the MPI wrapper KaMPIng [26].
@@ 3.5 Error handling
For invariants that can only be verified during execution, we use a layered assertion system: We verify invariants ranging from lightweight checks to assertions involving additional communication. These can be disabled level-by-level at compile-time, encouraging developers to use exhaustive checking while writing and test code, yet permitting a lean, high-performance configuration for production builds. By integrating error-handling controls directly into the C++ interface—rather than relying on rebuilding the MPI library itself in a “debug” mode—we provide an ergonomic, flexible framework.
Errors that may be recoverable are signaled by throwing exceptions. Here, special care only needs to be taken when throwing exceptions from destructors, since they may result in unexpected behavior. C++-23’s std::expected<T, E> offers a lightweight, value-based alternative to exceptions by encapsulating either a valid T or an error E in one object. A return-by-value design (see Section 3.4) fits naturally with the non-throwing, monadic error-propagation style.
@@ 4 Considerations Beyond a Preliminary C++ Interface
In the following, we discuss additional features which improve the usability of preliminary C++ MPI language bindings. These features are perhaps beyond the scope of initial standardization, since they introduce additional overhead on top of the current MPI specification, such as allocation or communication under the hood. Nevertheless, they make working with MPI from C++ easier and often safer. In the following, we discuss how to easily communicate data not directly compatible with MPI (Section 4.1), how to improve the handling of type lifetimes (Section 4.2) and to improve programmer productivity and reduce boilerplate code by offering sane defaults to programmers (Section 4.3).
@@ 4.1 Serialization
The data buffer concept defined in Section 3.3 provides a type-safe and easy-to-use way to use standard C++ containers and user-defined types with a regular memory layout and a direct mapping to an MPI data type in MPI communication. Commonly used complex C++ types, such as std::list, std::unordered_map, or polymorphic class hierarchies, do not provide contiguous storage or a well-defined memory representation. These are dynamic types: their size or layout may vary at runtime, and they may require logic to (re)construct from memory (on both ends of the communication) and may involve intermediate allocations associated with data copies. While such types cannot be communicated directly using the MPI type system, a user-defiend data buffer encapsulating packed data data can be used.But constructing this is cumbersome and error prone.
@@
Importantly, serialization libraries such as Boost.Serialization [20] or Cereal [12] let users define what data members of a type should be serialized; that is, what constitutes the object’s logical state. The concrete binary layout or packing is handled automatically by the library and is treated as an implementation detail  [^4] Implementations may even avoid byte packing and communicate in chunks.. This cleanly separates the intent of serialization from its mechanics, which is especially useful for complex types. Constructing such a serialized buffer may itself involve MPI communication, for example to exchange message sizes ahead of receiving the actual payload. While this is beyond the scope in terms of initial standardization, it can be built naturally on top of the previously defined data buffer concept. A user-defined type wrapping the serialization logic and exposing the required interface (contiguous, sized, and typed buffer) integrates seamlessly with the rest of the system.
@@ 4.2 Safer usage of complex types
The type classification and data buffer concept introduced in Sections 3.2 and 3.3 provide all necessary building blocks to work with complex types. Even if a type is not directly supported by MPI, for the large class of trivially copyable types, there still exists a type-safe, direct type mapping. This allows the user to semi-automatically construct a type, and pass it alongside a container to the communication call as a data buffer, as shown in Fig. 5.
But this has shortcomings: Committing an MPI data type is not free, since implementations may optimize the internal representation of a type when it is committed. So in a C++ interface, it is the user’s responsibility to commit the type and free it when it is not longer needed (by dropping ownership using the RAII pattern). This introduces additional problems, since now the user has to keep the type around all the time, and the type correspondence between the container and the data type object is lost. To improve usability here, we suggest the idea of a type pool that allows to store committed data types and to perform a lookup based on type information.
@@
There is one caveat to this approach. The lifetime of types in MPI is currently not clearly defined: In the session model, the lifetime of types is allowed to span across multiple sessions which suggests that the MPI type system consists orthogonally to MPI initialization. But it also mandates that types may only be created once a session or the world model is initialized, yet users are encouraged to commit types separately for each session.
This inconsistency appears to be unintentional but has not been resolved since sessions were introduced [^5] https://github.com/mpi-forum/mpi-issues/issues/733. Fixing this will enable language bindings to MPI, not only limited to C++, to handle types more easily. For example, when type lifetime is tied to sessions, a C++ implementation can associate a session-local type pool with each session object, which can queried for types without providing a pool explicitly. Types can then be freed automatically when the session is finalized.
@@ 4.3 Improving MPI productivity with sensible defaults
Many libraries improve usability features beyond a high-level standardized interface: automatically resizing receive containers, computing missing counts or displacements irregular collectives, or even performing additional communication to determine missing arguments. For example, Boost.MPI [13] and RWTH-MPI [9] offer various overloads for communication calls with default arguments, prevent users writing a lot of boilerplate code.
@@ 5 Conclusions
While many desire a truly modern C++ interface for MPI, this paper has laid out the conceptual and technical challenges needed to address before such interface can be standardized. Our goal was to clarify these challenges and provide a foundation of design principles that could guide future development and community discussion. To this end, we covered high-level considerations: a C++representation of the MPI object model (Section 3.1), data representation (Section 3.3) and ownership (Section 3.4), modeling and mapping types (Section 3.2), and idiomatic error handling (Section 3.5). We also discuss ideas beyond these core concepts, including serialization (Section 4.1) and usability improvements (Sections 4.2 and 4.3).
Based on these considerations, we identified specific aspects where the MPI standard requires clarification—most notably the lifecycle of user-defined data types, which remains underspecified in the context of sessions and makes automated resource management more difficult. We also support ongoing efforts to improve attribute support across the standard, particularly for sessions and requests, as this is essential for associating higher-level interface constructs with the lifetime of MPI objects (see Issue #664).
Finally, while callback support already exists in the standard, not all callbacks currently support user-defined state. Enabling this consistently would improve the design and usability of language bindings (see Issue #839). Standardizing a modern C++ interface is an ambitious – and perhaps impractical – undertaking. Instead, we advocate for the creation of language support guideline documents, similar in spirit to what this work aims to contribute.
We hope that our discussion can serve as a blueprint for other MPI language communities. A prime example for language support is Vapaa [14], a standalone implementation of MPI’s Fortran interface built on the C API, which strongly influenced the addition of an MPI Application Binary Interface in the upcoming MPI 5.0 standard [15]. Also, many of the mentioned concepts can be applied to or were inspired from the Rust MPI bindings rsmpi [23], since modern C++ and Rust share many conceptual similarities. For other languages, it might be helpful to derive their own, independent set of guidelines. This will allow MPI continue to evolve together with popular languages that define the future of high-performance computing.
@@ Acknowledgements
PSAAP Funding in part is acknowledged from these NSF Grants CCF-2405142 and CCF-2412182, as well as OAC-2103510 and the U.S. Department of Energy’s National Nuclear Security Administration (NNSA) under the Predictive Science Academic Alliance Program (PSAAP-III), Award DE-NA0003966. AAC work performed under the auspices of the US Department of Energy by Lawrence Livermore National Laboratory under contract DE-AC52-07NA27344,

### v2

@@ 3.4 Ownership and Non-Blocking Communication
Using the data buffer concept instead of raw pointers already offers benefits such as an ergonomic interaction with standard library containers, improved type safety and the potential to prevent out-of-bounds accesses. Another important advantage of this object-oriented data handling is that it allows us to model a proper concept of ownership. In C++, if an object owns a resource—such as heap-allocated memory—the object is responsible for releasing this resource when it goes out of scope. This is usually achieved by placing the clean-up code in the object’s destructor and the key idea behind the RAII idiom. A data buffer owning its underlying memory is therefore responsible for freeing this memory once it goes out of scope, thus preventing memory leaks. This plays well with C++’s move semantics. During move-construction or move-assignment, ownership of the underlying resource is transferred from the source to the destination object. Using this mechanism it is for example possible to move a (receive) data buffer to an MPI call as shown in Fig.˜6. Then the MPI call receives data into the provided buffer, which it now owns. Once the communication is completed, it returns the updated receive buffer by value resulting in a clean and idiomatic pass through of the underlying memory resource. This data flow allows us to avoid the use of out parameters that are discouraged in C++ [26, F.20].
In terms of library interface design, this “pass through” of data can be achieved by considering the value category of data buffers passed to communication calls. We propose that a C++ interface should explicitly distinguish between arguments passed as r- or l-values: If a data buffer is an r-value, the data buffer is moved to the MPI operation, and returned to the caller as a return value of the function call. If a data buffer is an l-value, it is passed by reference to the operation, ownership is not transferred, and it is not returned. For an example see Fig.˜6.
@@
A robust solution to this problem is to transfer ownership of data buffers involved in non-blocking communication calls to the C++ interface, which forward it to request object as show in Fig.˜6. The request object is conceptually similar to std::future  [^3] Using std::future to provide a safe interface for non-blocking communicating is not possible, as they are tied to asynchronous progress happening in the background, which the MPI standard does not guarantee.: Users can only access the data upon completing the request via wait or test, which move the data back to the caller, either directly by value in the former case, or encapsulated in std::optional in the latter. This way, invalid accesses to data involved in non-blocking communication can be prevented entirely through library semantics. This idea was first introduced and implemented in the MPI wrapper KaMPIng [27].
@@ 3.5 Error Handling
For invariants that can only be verified during execution, we use a layered assertion system: We verify invariants ranging from lightweight checks to assertions involving additional communication. These can be disabled level-by-level at compile-time, encouraging developers to use exhaustive checking while writing and testing code, yet permitting a lean, high-performance configuration for production builds. By integrating error-handling controls directly into the C++ interface—rather than relying on rebuilding the MPI library itself in a “debug” mode—we provide an ergonomic, flexible framework. If an assertion is enabled and fails, we call MPI_Abort because execution can not safely continue.
Errors that may be recoverable are signaled to the user. This can either by done by throwing an exception, or by returning std::expected from MPI procedures. C++-23’s std::expected<T, E> offers a lightweight, value-based alternative to exceptions, by encapsulating either a valid T or an error E in one object. Using std::expected makes error handling explicit, avoids hidden blocking operations inside distant catch handlers, and allows for easy error chaining and propagation. Together with the return-by-value-based design proposed in Section˜3.4, this extends the concept of error codes in MPI’s C interface, by only providing access to the return value of a procedure in case of success.
@@ 4 Considerations Beyond a Preliminary C++ Interface
In the following, we discuss additional features which improve the usability of preliminary C++ MPI language bindings. These features are perhaps beyond the scope of initial standardization, since they introduce additional overhead on top of the current MPI specification, such as allocation or communication under the hood. Nevertheless, they make working with MPI from C++ easier and often safer. In the following, we discuss how to easily communicate data not directly compatible with MPI (Section˜4.1), how to improve the handling of type lifetimes (Section˜4.2) and to improve programmer productivity and reduce boilerplate code by offering sane defaults to programmers (Section˜4.3).
@@ 4.1 Serialization
The data buffer concept defined in Section˜3.3 provides a type-safe and easy-to-use way to use standard C++ containers and user-defined types with a regular memory layout and a direct mapping to an MPI data type in MPI communication. Commonly used complex C++ types, such as std::list, std::unordered_map, or polymorphic class hierarchies, do not provide contiguous storage or a well-defined memory representation. These are dynamic types: their size or layout may vary at runtime, they may require logic to (re)construct from memory (on both ends of the communication) and may involve intermediate allocations associated with data copies. While such types cannot be communicated directly using the MPI type system, a user-defined data buffer encapsulating packed data data can be used. However, constructing this is cumbersome and error-prone.
@@
Importantly, serialization libraries such as Boost.Serialization [21] or Cereal [12] let users define what data members of a type should be serialized; that is, what constitutes the object’s logical state. The concrete binary layout or packing is handled automatically by the library and is treated as an implementation detail  [^4] Implementations may even avoid byte packing and communicate in chunks.. This cleanly separates the intent of serialization from its mechanics, which is especially useful for complex types. Constructing such a serialized buffer may itself involve MPI communication, for example to exchange message sizes ahead of receiving the actual payload. While this is beyond the scope of the MPI standard, it can be built naturally on top of the previously defined data buffer concept. A user-defined type wrapping the serialization logic and exposing the required interface (contiguous, sized, and typed buffer) integrates seamlessly with the rest of the system.
@@ 4.2 Safer Usage of Complex Types
The type classification and data buffer concept introduced in Sections˜3.2 and 3.3 provide all necessary building blocks to work with complex types. Even if a type is not directly supported by MPI, for the large class of trivially copyable types, there still exists a type-safe, direct type mapping. This allows the user to semi-automatically construct a type, and pass it alongside a container to the communication call as a data buffer, as shown in Fig.˜5.
But this has shortcomings: Committing an MPI data type is not free, since implementations may optimize the internal representation of a type when it is committed. So in a C++ interface, it is the user’s responsibility to commit the type and free it when it is no longer needed (by destructing the type object using the RAII pattern). This introduces additional problems, since now the user has to keep the type around all the time, and the type correspondence between the container and the data type object is lost. To improve usability here, we suggest the idea of a type pool that allows to store committed data types and to perform a lookup based on type information.
@@
There is one caveat to this approach: The lifetime of types in MPI is currently not clearly defined. In the session model, the lifetime of types is allowed to span across multiple sessions which suggests that the MPI type system exists orthogonally to MPI initialization. But it also mandates that types may only be created once a session or the world model is initialized, yet users are encouraged to commit types separately for each session.
This inconsistency appears to be unintentional but has not been resolved since sessions were introduced [^5] https://github.com/mpi-forum/mpi-issues/issues/733. Fixing this will enable language bindings to MPI, not only limited to C++, to handle types more easily. For example, when type lifetime is tied to sessions, a C++ implementation can associate a session-local type pool with each session object, which can be queried for types without providing a pool explicitly. Types can then be freed automatically when the session is finalized.
@@ 4.3 Improving MPI Productivity with Sensible Defaults
Many libraries improve usability features beyond a high-level standardized interface: automatically resizing receive containers, computing missing counts or displacements in collectives with varying counts, or even performing additional communication to determine missing arguments. For example, Boost.MPI [13] and RWTH-MPI [9] offer various overloads for communication calls with default arguments, preventing users from writing a lot of boilerplate code.
@@ 5 Conclusions
Standardizing a modern C++ interface is an ambitious—and perhaps impractical—undertaking. Instead, we advocated here for the creation of language support guideline documents, similar in spirit to what this work aims to contribute.
While many desire a truly modern C++ interface for MPI, this paper has laid out the conceptual and technical challenges that need to be addressed before such an interface can be standardized. Our goal was to clarify these challenges and provide a foundation of design principles that could guide future development and community discussion. To this end, we covered high-level considerations: a C++representation of the MPI object model (Section˜3.1), data representation (Section˜3.3) and ownership (Section˜3.4), modeling and mapping types (Section˜3.2), and idiomatic error handling (Section˜3.5). We also discussed ideas beyond these core concepts, including serialization (Section˜4.1) and usability improvements (Sections˜4.2 and 4.3).
Based on these considerations, we identified specific aspects where the MPI standard requires clarification—most notably the lifecycle of user-defined data types, which remains under-specified in the context of sessions and makes automated resource management more difficult. We also support ongoing efforts to improve attribute support across the standard, particularly for sessions and requests, as this is essential for associating higher-level interface constructs with the lifetime of MPI objects (see Issue #664). Finally, while callback support already exists in the standard, not all callbacks currently support user-defined state. Enabling this consistently would improve the design and usability of language bindings (see Issue #839).
We hope that our discussion can serve as a blueprint for other MPI language communities. A prime example for language support is Vapaa [14], a standalone implementation of MPI’s Fortran interface built on the C API, which strongly influenced the addition of an MPI Application Binary Interface in the MPI 5.0 standard [15]. Also, many of the mentioned concepts can be applied to or were inspired from the Rust MPI bindings rsmpi [24], since modern C++ and Rust share many conceptual similarities. For other languages, it might be helpful to derive their own, independent set of guidelines. This will enable MPI continue to evolve together with popular languages that define the future of high-performance computing.
{credits}
@@ Acknowledgements
PSAAP Funding in part is acknowledged from these NSF Grants OAC-2514054, CNS-2450093, CCF-2405142, and CCF-2412182 and the U.S. Department of Energy’s National Nuclear Security Administration (NNSA) under the Predictive Science Academic Alliance Program (PSAAP-III), Award DE-NA0003966. AAC work performed under the auspices of the US Department of Energy by Lawrence Livermore National Laboratory under contract DE-AC52-07NA27344, and supported by the Center for Non-Perturbative Studies of Functional Materials Under Non-Equilibrium Conditions (NPNEQ) funded by the Computational Materials Sciences Program of the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. This work was also performed under the auspices of the US Department of Energy’s Pacific Northwest National Laboratory, operated by Battelle Memorial Institute under contract DE-AC05-76RL01830. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation, or the U.S. Department of Energy’s National Nuclear Security Administration.